# A uniform-in-time nonlocal approximation of the standard Fokker-Planck equation

### v2

@@ 3.3 An $L^{\infty}$ version of Berry-Esseen’s central limit theorem
The first two terms can be bounded in a similar way. For $T_{1}$, and without loss of generality, assume $p^{\prime}$ is an integer (otherwise we can work with $\left\lceil p^{\prime}\right\rceil$, the smallest integer larger or equal than $p^{\prime}$, since $\hat{f}$ is also in $L^{\left\lceil p^{\prime}\right\rceil}$). Then
$\int_{B_{\delta}^{c}}\prod_{i=1}^{n}\left|\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|\,\mathrm{d}\xi=n^{d/2}\int_{|\eta|\geq\frac{L}{l}\delta}\prod_{i=1}^{n}\left|\hat{f}\Big{(}\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big{)}\right|\,\mathrm{d}\eta\\ \leq n^{d/2}\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big{(}\frac{\sigma_{p^{\prime}+1}}{\bar{\sigma}}\eta\Big{)}\right|\cdots\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big{(}\frac{\sigma_{n}}{\bar{\sigma}}\eta\Big{)}\right|\int_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big{(}\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big{)}\right|\cdots\left|\hat{f}\Big{(}\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big{)}\right|\,\mathrm{d}\eta.$ (18)
For each $i=p^{\prime}+1,\dots,n$
$\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big{(}\frac{\sigma_{i}}{\bar{\sigma}}\eta\Big{)}\right|\leq\sup_{|\eta|\geq\frac{\bar{\sigma}}{\sigma_{i}}\delta}\left|\hat{f}\Big{(}\frac{\sigma_{i}}{\bar{\sigma}}\eta\Big{)}\right|=\sup_{|\eta^{\prime}|\geq\delta}|\hat{f}(\eta^{\prime})|\leq\kappa(\delta)$ (19)
with $\kappa(\delta)\in(0,1)$, coming from Lemma 3.9 $(i)$. The factors inside the integral in Equation (18) can be bounded by using Hölder’s inequality: we rewrite
$\begin{split}\int_{|\eta|\geq\frac{L}{l}\delta}&\left|\hat{f}\Big{(}\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big{)}\right|\cdots\left|\hat{f}\Big{(}\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big{)}\right|\,\mathrm{d}\eta\\ &\leq\bigg{(}\int_{\mathbb{R}^{d}}\left|\hat{f}\Big{(}\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big{)}\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg{)}^{1/p^{\prime}}\cdots\bigg{(}\int_{\mathbb{R}^{d}}\left|\hat{f}\Big{(}\frac{\sigma_{n}}{\bar{\sigma}}\eta\Big{)}\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg{)}^{1/p^{\prime}}\\ &=\Big{(}\frac{\sigma_{1}}{\bar{\sigma}}\Big{)}^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\cdots\Big{(}\frac{\sigma_{n}}{\bar{\sigma}}\Big{)}^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\leq\Big{(}\frac{L}{l}\Big{)}^{d}\|\hat{f}\|_{p^{\prime}}^{p^{\prime}}\leq\Big{(}\frac{L}{l}\Big{)}^{d}C_{p}\|f\|_{p}^{p^{\prime}},\end{split}$ (20)
@@
Proceed similarly for $T_{2}$ to get that there exists $C_{1}$ such that
$T_{1}+T_{2}\leq\frac{C_{1}}{n^{s/2}}\qquad\text{for $n>p^{\prime}$.}$
Now we focus our attention on $T_{3}$. Using the factorisation in Lemma 3.7 we obtain
$\left|\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}\right|=\prod_{i=1}^{n}\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}-\prod_{i=1}^{n}\hat{G}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}\\ =\sum_{i=1}^{n}\left|\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}-\hat{G}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|\left|\prod_{j=1}^{i-1}\hat{f}\Big{(}\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|\left|\prod_{j=i+1}^{n}\hat{G}\Big{(}\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|.$
@@
$\left|\hat{f}\Big{(}\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|\leq e^{-\frac{|\xi|^{2}}{4n}\frac{\sigma_{i}^{2}}{\bar{\sigma}^{2}}}\leq e^{-\frac{|\xi|^{2}}{4n}(\frac{l}{L})^{2}}.$
@@
$\left|\hat{G}\Big{(}{\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}}\Big{)}\right|=(2\pi)^{-d/2}e^{-\frac{|\xi|^{2}}{2n}\frac{\sigma_{j}^{2}}{\bar{\sigma}^{2}}}\leq e^{-\frac{|\xi|^{2}}{2n}(\frac{l}{L})^{2}}.$
@@
$\begin{split}\frac{\left|\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}\right|}{|\xi|^{2+s}}&=\frac{1}{\left|\xi\right|^{2+s}}\Big{[}\sum_{i=1}^{n}\left|\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}-\hat{G}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|\left|\prod_{j=1}^{i-1}\hat{f}\Big{(}\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|\left|\prod_{j=i+1}^{n}\hat{G}\Big{(}\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|\Big{]}\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{i=1}^{n}\left|\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}-\hat{G}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|e^{-\frac{|\xi|^{2}}{4n}(i-1)(\frac{l}{L})^{2}}e^{-\frac{|\xi|^{2}}{2n}(n-1)(\frac{l}{L})^{2}}\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{i=1}^{n}\left|\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}\frac{\left|\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}-\hat{G}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}\right|}{\left|\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}}e^{-\frac{|\xi|^{2}}{4}\frac{n-1}{n}(\frac{l}{L})^{2}}\\ &\leq\frac{1}{n^{\frac{2+s}{2}}}\Big{(}\frac{L}{l}\Big{)}^{2+s}n\sup_{\eta^{\prime}}\frac{|\hat{f}(\eta^{\prime})-\hat{G}(\eta^{\prime})|}{\left|\eta^{\prime}\right|^{2+s}}e^{-\frac{\left|\xi\right|^{2}}{8}(\frac{l}{L})^{2}}.\end{split}$
Since $f$ and $G$ are both centred and have the same covariance matrix, in the same spirit of the proof of Lemma 3.9 $(i)$, we obtain
@@
$\begin{split}T_{3}&\leq\frac{C_{1}}{n^{s/2}}\Big{(}\frac{L}{l}\Big{)}^{2+s}(\rho_{2+s}+M_{2+s}(G))\int_{\mathbb{R}^{d}}e^{-\frac{\left|\xi\right|^{2}}{8}(\frac{l}{L})^{2}}\,\mathrm{d}\xi\\ &:=\frac{1}{n^{s/2}}\Big{(}\frac{L}{l}\Big{)}^{2+s}(\rho_{2+s}+M_{2+s}(G))C_{\frac{L}{l},d}.\end{split}$
Adding up $T_{1},T_{2}$ and $T_{3}$, the proof is completed. ∎
@@ 3.4 Positivity of the solution
Let $X_{m}$ be a Poisson random variable of parameter $m$. For a fixed $m$, using a known result on bounds for the tail probabilities of a Poisson random variable, we have that
$s_{m}=1-\mathbb{P}(X_{m}<m)-\mathbb{P}(X_{m}\leq 2m+1)\geq\frac{1}{2}-e^{m+1}\Big{(}\frac{m}{2m+1}\Big{)}^{2m+1}:=b_{m}\geq b_{1}=:C_{L},$
@@
Expressing the solution via Wild sums we have.
$u(t,x)\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}(e^{t}x)\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}$
@@
$\begin{split}u(t,x)&\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{\mathbb{R}^{d}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{B_{R_{2}}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq Ae^{dt}e^{-\frac{t}{\varepsilon^{2}}}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\Big{(}\frac{t}{\varepsilon^{2}}\Big{)}^{n}\frac{1}{n!}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\\ &\geq\Big{(}Ae^{\,\mathrm{d}t}e^{-(\frac{t}{\varepsilon^{2}}-\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor)}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\Big{)}e^{-\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor^{n}\frac{1}{n!}\\ &\geq\Big{(}Ae^{dt-1}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\Big{)}e^{-m}\sum_{n=m}^{2n}\frac{m^{n}}{n!}\geq AC_{L}e^{dt-1}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\end{split}$
@@
$\begin{split}u(t,x)&\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{\mathbb{R}^{d}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{B_{R_{2}}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=N}^{\infty}\Big{(}\frac{t}{\varepsilon^{2}}\Big{)}^{n}\frac{1}{n!}B_{n}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\\ &\geq\frac{B}{N!}e^{(d-\frac{1}{\tilde{\varepsilon}_{0}^{2}})t}\Big{(}\frac{t}{\varepsilon_{0}^{2}}\Big{)}^{N}\frac{1}{N^{\frac{d}{2}}}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y.\end{split}$
@@
$\alpha:=\min\Big{\{}AC_{L}e^{dt-1},\frac{B}{N!}e^{(d-\frac{1}{\varepsilon_{0}^{2}})t}\Big{(}\frac{t}{\tilde{\varepsilon}_{0}^{2}}\Big{)}^{N}\frac{1}{N^{\frac{d}{2}}}\Big{\}}.\qed$
@@
Theorem 3.2 gives the confining condition on the generator $L_{\varepsilon}$ in $L^{1}_{\left\langle k\right\rangle}$ and consequently in $L^{1}_{k}$. Let $u_{0}$ be in $D(L_{\varepsilon})$. By standard semigroup theory $u(t)\in\mathcal{C}^{1}([0,\infty);L^{1}_{k})$, and consequently we compute with $\varphi=\left\langle x\right\rangle^{k}$,
@@
where to prove $*\stackrel{{\scriptstyle\mathclap{\mbox{*}}}}{{=}}$ we use a classical cutoff argument with a bump function. We then use Gronwall’s lemma with the $\mathcal{C}^{1}$ function $f(t):=\int_{\mathbb{R}^{d}}u(t)\varphi$: That is,
@@
Since $D(L)$ is dense in $L^{1}_{k}$, the claim is true also for any $u_{0}\in L^{1}_{k}$ by standard density arguments, hence proving the Lyapunov condition on semigroups. Similarly, part $(ii)$ of Theorem 1.1 is an immediate consequence of Theorem 3.10 and Theorem 3.4, proving a semigroup Lyapunov condition in a similar manner with $\varphi=e^{a\left\langle x\right\rangle}.$

### v3

@@ 3.3 An $L^{\infty}$ version of Berry-Esseen’s central limit theorem
The first two terms can be bounded in a similar way. For $T_{1}$, and without loss of generality, assume $p^{\prime}$ is an integer (otherwise, we can work with $\left\lceil p^{\prime}\right\rceil$, the smallest integer larger or equal than $p^{\prime}$, since $\hat{f}$ is also in $L^{\left\lceil p^{\prime}\right\rceil}$). Then
$\int_{B_{\delta}^{c}}\prod_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\,\mathrm{d}\xi=n^{d/2}\int_{|\eta|\geq\frac{L}{l}\delta}\prod_{j=1}^{n}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta\\ \leq n^{d/2}\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}+1}}{\bar{\sigma}}\eta\Big)\right|\cdots\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{n}}{\bar{\sigma}}\eta\Big)\right|\int_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\right|\cdots\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta.$ (18)
For each $j=p^{\prime}+1,\dots,n$
$\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|\leq\sup_{|\eta|\geq\frac{\bar{\sigma}}{\sigma_{i}}\delta}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|=\sup_{|\eta^{\prime}|\geq\delta}|\hat{f}(\eta^{\prime})|\leq\kappa(\delta)$ (19)
with $\kappa(\delta)\in(0,1)$, coming from Lemma 3.9 $(ii)$. The factors inside the integral in Equation (18) can be bounded by using Hölder’s inequality: we rewrite
$\begin{split}\int_{|\eta|\geq\frac{L}{l}\delta}&\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\right|\cdots\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta\\ &\leq\bigg(\int_{\mathbb{R}^{d}}\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg)^{1/p^{\prime}}\cdots\bigg(\int_{\mathbb{R}^{d}}\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg)^{1/p^{\prime}}\\ &=\Big(\frac{\sigma_{1}}{\bar{\sigma}}\Big)^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\cdots\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\Big)^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\leq\Big(\frac{L}{l}\Big)^{d}\|\hat{f}\|_{p^{\prime}}^{p^{\prime}}\leq\Big(\frac{L}{l}\Big)^{d}C_{p}\|f\|_{p}^{p^{\prime}},\end{split}$ (20)
@@
We proceed analogously for $T_{2}$, obtaining a very similar bound also for $T_{2}$. Since $\kappa(\delta)<1$, $\kappa(\delta)^{n-p^{\prime}}$ decays very fast in $n$, thus for $n$ large enough (depending on $p$), there exists $C_{1}>0$ such that
$T_{1}+T_{2}\leq\frac{C_{1}}{n^{s/2}}$ (22)
Now we focus our attention on $T_{3}$. Using the factorization in Lemma 3.7 we obtain
$\left|\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi)\right|=\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\prod_{i=1}^{n}\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\\ =\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=1}^{j-1}\hat{f}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=j+1}^{n}\hat{G}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|.$
@@
$\left|\hat{f}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\leq e^{-\frac{|\xi|^{2}}{4n}\frac{\sigma_{k}^{2}}{\bar{\sigma}^{2}}}\leq e^{-\frac{|\xi|^{2}}{4n}(\frac{l}{L})^{2}}.$
@@
$\left|\hat{G}\Big({\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}}\Big)\right|=(2\pi)^{-d/2}e^{-\frac{|\xi|^{2}}{2n}\frac{\sigma_{k}^{2}}{\bar{\sigma}^{2}}}\leq e^{-\frac{|\xi|^{2}}{2n}(\frac{l}{L})^{2}}.$
@@
$\begin{split}\frac{\left|\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi)\right|}{|\xi|^{2+s}}&=\frac{1}{\left|\xi\right|^{2+s}}\Big[\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=1}^{j-1}\hat{f}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=j+1}^{n}\hat{G}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\Big]\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|e^{-\frac{|\xi|^{2}}{4n}(j-1)(\frac{l}{L})^{2}}e^{-\frac{|\xi|^{2}}{2n}(n-j)(\frac{l}{L})^{2}}\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{j=1}^{n}\left|\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}\frac{\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|}{\left|\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}}e^{-\frac{|\xi|^{2}}{4}\frac{n-1}{n}(\frac{l}{L})^{2}}\\ &\leq\frac{1}{n^{\frac{2+s}{2}}}\Big(\frac{L}{l}\Big)^{2+s}n\sup_{\eta^{\prime}}\frac{|\hat{f}(\eta^{\prime})-\hat{G}(\eta^{\prime})|}{\left|\eta^{\prime}\right|^{2+s}}e^{-\frac{\left|\xi\right|^{2}}{8}(\frac{l}{L})^{2}}.\end{split}$
Since $f$ and $G$ are both centered and have the same covariance matrix, in the same spirit of the proof of Lemma 3.9 $(i)$, we obtain
@@
$\begin{split}T_{3}&\leq\frac{C_{1}}{n^{s/2}}\Big(\frac{L}{l}\Big)^{2+s}(\rho_{2+s}+M_{2+s}(G))\int_{\mathbb{R}^{d}}e^{-\frac{\left|\xi\right|^{2}}{8}(\frac{l}{L})^{2}}|\xi|^{2+s}\,\mathrm{d}\xi\\ &:=\frac{1}{n^{s/2}}\Big(\frac{L}{l}\Big)^{2+s}(\rho_{2+s}+M_{2+s}(G))C_{\frac{L}{l},d}.\end{split}$
Taking into account (22) and adding $T_{3}$, the proof is completed. ∎
@@ 3.4 Positivity of the solution
the la (not the updated version) Let $X_{m}$ be a Poisson random variable of parameter $m$. For a fixed $m$, using a known result on bounds for the tail probabilities of a Poisson random variable, we have that
$s_{m}=1-\P(X_{m}<m)-\P(X_{m}\leq 2m+1)\geq\frac{1}{2}-e^{m+1}\Big(\frac{m}{2m+1}\Big)^{2m+1}:=b_{m}\geq b_{1}=:C_{L},$
@@
Expressing the solution via Wild sums we have
$u(t,x)\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}(e^{t}x)\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}.$
@@
$\begin{split}u(t,x)&\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{\mathbb{R}^{d}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{B_{R_{2}}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq Ae^{dt}e^{-\frac{t}{\varepsilon^{2}}}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\Big(\frac{t}{\varepsilon^{2}}\Big)^{n}\frac{1}{n!}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\\ &\geq\Big(Ae^{\,\mathrm{d}t}e^{-(\frac{t}{\varepsilon^{2}}-\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor)}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\Big)e^{-\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor^{n}\frac{1}{n!}\\ &\geq\Big(Ae^{dt-1}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\Big)e^{-m}\sum_{n=m}^{2m}\frac{m^{n}}{n!}\geq AC_{L}e^{dt-1}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\end{split}$
@@
$\begin{split}u(t,x)&\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{\mathbb{R}^{d}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{B_{R_{2}}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=N}^{\infty}\Big(\frac{t}{\varepsilon^{2}}\Big)^{n}\frac{1}{n!}B_{n}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\\ &\geq\frac{B}{N!}e^{(d-\frac{1}{\tilde{\varepsilon}_{0}^{2}})t}\Big(\frac{t}{\varepsilon_{0}^{2}}\Big)^{N}\frac{1}{N^{\frac{d}{2}}}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y.\end{split}$
@@
$\alpha:=\min\Big\{AC_{L}e^{dt-1},\frac{B}{N!}e^{(d-\frac{1}{\varepsilon_{0}^{2}})t}\Big(\frac{t}{\tilde{\varepsilon}_{0}^{2}}\Big)^{N}\frac{1}{N^{\frac{d}{2}}}\Big\}.$
@@
(put it twice: once after Lyapunov and once here) Theorem 3.2 gives the confining condition on the generator $L_{\varepsilon}$ in $L^{1}_{\left\langle k\right\rangle}$ and consequently in $L^{1}_{k}$. Let $u_{0}$ be in $\mathcal{D}(L_{\varepsilon})$. By standard semigroup theory $u(t)\in\mathcal{C}^{1}([0,\infty);L^{1}_{k})$, and consequently we compute with $\varphi=\left\langle x\right\rangle^{k}$,
@@
[…]
@@
Since $\mathcal{D}(L_{\varepsilon})$ is dense in $L^{1}_{k}$, the claim is true also for any $u_{0}\in L^{1}_{k}$ by standard density arguments, hence proving the Lyapunov condition on semigroups. Similarly, part $(ii)$ of Theorem 1.1 is an immediate consequence of Theorem 3.10 and Theorem 3.4, proving a semigroup Lyapunov condition in a similar manner with $\varphi=e^{a\left\langle x\right\rangle}.$
Remark 3.16 (On the dependence of the constants).
As mentioned in the introduction, Harris’s theorem provides a constructive method that yields explicit—though generally nonoptimal—constants. Unsurprisingly, the kernel $J$ plays a central role in determining these constants. Indeed, the latter get better the closer $J$ resembles a Gaussian distribution. This is reflected in our use of the Berry–Esseen theorem, which introduces an explicit dependence on both the Berry–Esseen constant and the number $N$ of convolution iterations. We also note that if $J$ were itself a Gaussian, positivity would follow directly and there would be no need to invoke any central limit theorem.